Do It! Review 11-3b Crane Company has had 4 years of record earnings. Due to this success, the market price of its 385,000 shares of $2 par value common stock has increased from $13 per share to $53. During this period, paid-in capital remained the same at $2,310,000. Retained earnings increased from $1,732,500 to $11,550,000. CEO Don Ames is considering either (1) a 15% stock dividend or (2) a 2-for-1 stock split. He asks you to show the before-and-after effects of each option on (a) retained earnings, (b) total stockholders equity, and (c) par value per share 1. Stock dividend - retained earnings 2. 2-for-1 stock split retained earnings Crane Company Original Balance After DividendAfter Split Paid-in capital Retained earnings Total stockholders equity s Shares outstanding 1. Stock dividend - par value per share 2. 2-for-1 stock split par value per share
Answer #1

a)

1) 15% stock dividend

385,000*15% = 57,750 share

Retained earnings =(57,750*53) = 3,060,750

2) stock dividend -retained earnings = 3,060,750

2-for -1 stock split- retained earnings = 0

b)

original balance after dividend after split
paid in capital 2,310,000 5,370,750(2,310,000+3,060,750) 2,310,000
retained earnings 11,550,000 8,489,250(11,550,000-3,060,750) 11,550,00
total stockholders Equity 13,860,000 13,860,000 13,860,000
shares outstanding 385,000 442,750[385000+57750] 770,000[385000*2]

c)

1)Stock dividend - par value per share=$2

2)2-for-1 stock split - par value per share = $2/2 =>$1